Your firm owns two plants that produce chocolate of exactly the same quality. One plant has a cost function given by cA (y) = 4y2 while the other plant has a cost function given by cB (y ) = y2 + y + 4. Given that you can optimally choose how to distribute the production of chocolate across these two plants, what is your firm's marginal cost function for producing chocolate?
Answer #1

Let output produced in plant 1 be y1 and in plant 2 be y2.

CA = 4y12

MC(A) = dCA/dy1 = 8y1

y1 = MC(A) / 8 = 0.125 x MC(A)

CB = y22 + y2 + 4

MC(B) = dCB/dy2 = 2y2 + 1

2y2 = MC(B) - 1

y2 = 0.5MC x (B) - 0.5

Setting MC(A) = MC(B) = MC:

y1 = 0.125 x MC

y2 = 0.5 x MC - 0.5

y = y1 + y2

y = 0.125 x MC + 0.5 x MC - 0.5

y = MC x (0.125 + 0.5) - 0.5

y = MC x 0.625 - 0.5

MC x 0.625 = y + 0.5

MC = (y + 0.5) / 0.625 = 1.6y + 0.8